OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? 6(3x + 9) - 3 A. 18x + 6 B. 18x + 54 C. 18x + 51 D. 18x + 36

OpenStudy (azureilai):

Use the distributive property 6*3= 18 6*9=54 54-3=51 18x+51

OpenStudy (anonymous):

......ooooooo

OpenStudy (anonymous):

thx